Shift report 20020107 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon Jan 07 2002 - 00:01:27 EST

  • Next message: Apache: "Shift report 20020107 00:00-08:00"

    ------------------------------------------------------------
    Date: Mon Jan  7 00:01:27 2002, Shift: 16:00-24:00
    Supervisor: kh, with: 
    Log book#12 pages 9 - 13
    
    Shift Summary:
    ------------------------------------------------------------
    A fairly productive shift.  Had beam right at the beginning
    of the shift and started a 0 field run.  Shortly thereafter
    MCR called and asked if we wanted to keep the beam as it
    was not polarized.  I said keep it.  They called 5 minutes
    later saying they would dump because 4 out of the other 5
    experiments wanted to dump.  Came back about 18:00 and
    that has lasted until shift end.
    
    Achievements:
    Runs    trig 3     trig 6         FS                 MRS
           effective  effective
    6315     34908      12512    12 deg 0 field (D1=0)   35 4A
    6316     68881      22835    12 deg 0 field (D1=1/8) 35 4A
    6317     73264      23816    12 deg 1/8 field        35 4A
    6318    180525      63713    12 deg 1/8 field        40 4A
    
    Run 6318 still running at shift end.  Numbers indicate
    statistics at shift end.
    
    The run goal for MRS at 35 deg setting 4A was achieved with
    462k events effective.  I therefore moved the MRS to 40 
    as can be seen for run 6318
    
    Problems:
    1. T3 and T5 cathodes tripped.  Cause is unknown.  This
       was discovered quickly thanks to badtrip.
    
    2. Lost connection to Ioserver.  Regained by using
       instructions on "Share your Experience"  It worked.
    
    2. I noticed a water leak in the roof just above the 
       D5 power supply.  It had obviously been noticed before
       because there was black plastic guiding the water away
       from the power supply (more or less working)  But the
       bottom by the pumps had a fair amount of water.  Called
       MCR who sent support and they mopped it up and promised
       that someone would work on the roof tomorrow.
    
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Mon Jan 07 2002 - 00:02:20 EST